Solve the inequality. -x divided by 7 + 4 >= 3x
Answer by stanbon(75887)

Solve the inequality. -x divided by 7 + 4 >= 3x
------
-x/7 + 4 >= 3x
----
Add (x/7) to both sides to get:
------
(3 1/7)x <= 4
---
(22/7)x <= 4
x <= 4(7/22)
x <= 14/11
================
Cheers,
Stan H.